OpenStudy (anonymous):

...how do i do this?... Graph 3x + 9y = -21 and check on your calculator. What is the x-intercept, the y-intercept, and the slope?

OpenStudy (anonymous):

x-intercept is the x value when y equals to 0. Y-intercept is the y value when x equals to 0. the slope is -1/3

OpenStudy (anonymous):

So X =-7 and Y = -2.33?

OpenStudy (anonymous):

so basically x-int = - 7, y-int = -7/3, m = -1/3 no?

OpenStudy (anonymous):

 yep

OpenStudy (anonymous):

Alright cool Thanks Dude!

OpenStudy (lgbasallote):

small mistake..it;s not -1/3

OpenStudy (lgbasallote):

9y = -21 y = -21/9 y = -7/3